FORMALDEHYDE: THE 5.7μm AND 3.6μm BANDS

\underline{D. JACQUEMART}, N. LACOME, Université Pierre et Marie Curie-Paris 6; CNRS; Laboratoire de Dynamique, Interactions et Réactivité (LADIR), UMR 7075, Case Courrier 49, 4 Place Jussieu, 75252 Paris Cedex 05, France; A. PERRIN, F. KWABIA TCHANA, Université Paris 12; CNRS; Laboratoire Inter Universitaire des Systèmes Atmosphériques (LISA), UMR 7583, 61 Av du General de Gaulle, 94010 Créteil Cedex France; A. LARAIA, and R.R. GAMACHE, University of Mass Lowell, Department of Environmental, Earth and Atmospheric Sciences, Lowell, MA 01854, USA.

The goal of this study was to achieve absolute line intensities and reliable line broadening parameters for the strong 5.7 μm and 3.6 μm bands of formaldehyde (H_{2}CO) and to generate, for both spectral regions, a complete linelist for atmospheric applications.

High-resolution Fourier transform spectra were recorded at LADIR for the whole 1600 - 3200 cm^{−1} spectral range and for different path-length-pressure products conditions. Using these spectra, a large set of H_{2}CO individual line intensities^{a} and of self- and N_{2}-broadening linewidths^{b} were measured.

The calculated band intensities derived for the 5.7 μm and 3.6 μm bands are in excellent agreement with the values achieved recently by medium resolution band intensity measurements. It has to be mentioned that intensities in the 3.6 μm achieved in this work are, on the average, about 28 % stronger than those quoted in the HITRAN 2004 database. The linelist of positions and intensities based on this work^{a} has been included in the HITRAN 2008 database.

Concerning the linewidths, empirical expansions and theoretical calculations (semi-classical Robert-Bonamy formalism) have been performed and compared to the measurements. Rotational dependences have been studied, and the temperature dependence of the N_{2}-broadening coefficients has been calculated.

aA. Perrin, D. Jacquemart, F. Kwabia-Tchana, and N. Lacome, JQSRT 110,700-716,2009.

bD. Jacquemart, A. Laraia F. Kwabia-Tchana, R.R. Gamache A. Perrin and N. Lacome, JQSRT in press.